SUB-DOPPLER HIGH RESOLUTION EXCITATION SPECTROSCOPY OF DIBENZO-P-DIOXIN

\underline{MASAAKI BABA}, Graduate School of Science, Kyoto University, Kyoto 606-8501, Japan; ATSUSHI DOI, YOSHIO TATAMITANI, SHUNJI KASAHARA, and HAJIME KATO, Molecular Photoscience Research Center, Kobe University, Kobe 657-8501, Japan.

Sub-Doppler high-resolution fluorescence excitation spectra of the S_{1} ←S_{0} transition of dibenzo-p-dioxin (DD) have been observed in a collimated molecular beam.^{a} The rotational constants in both the S_{0} and S_{1} states were determined by analysis of the observed rotational contours. The DD molecule has shown to be planar in the S_{0} state and folded slightly out-of-plane (butterfly form) in the S_{1} state. The observed line widths were about 0.01 cm^{−1}, which were much larger than the instrumental resolution (0.001 cm^{−1}). The lifetime of the S_{1} state was evaluated to be about 500 ps.

aM. Baba, A. Doi, Y. Tatamitani, S. Kasahara, and H. Katô, J. Phys. Chem. A, 108, 1388 (2004)